# Single Image Rolling Shutter Removal with Diffusion Models

### v1

@@ Abstract
We present RS-Diffusion, the first Diffusion Models-based method for single-frame Rolling Shutter (RS) correction. RS artifacts compromise visual quality of frames due to the row-wise exposure of CMOS sensors. Most previous methods have focused on multi-frame approaches, using temporal information from consecutive frames for the motion rectification. However, few approaches address the more challenging but important single frame RS correction. In this work, we present an “image-to-motion” framework via diffusion techniques, with a designed patch-attention module. In addition, we present the RS-Real dataset, comprised of captured RS frames alongside their corresponding Global Shutter (GS) ground-truth pairs. The GS frames are corrected from the RS ones, guided by the corresponding Inertial Measurement Unit (IMU) gyroscope data acquired during capture. Experiments show that our RS-Diffusion surpasses previous single RS correction methods. Our method and proposed RS-Real dataset lay a solid foundation for advancing the field of RS correction.
Index Terms:
Diffusion models, homography, gyroscope, transformer.
I Introduction
@@
Rolling shutter (RS) is a common effect encountered when capturing images with CMOS sensors. It results from varying exposure times for different rows in each frame, causing artifacts like distorted straight lines and skewed image content, as shown in Fig. 1 (Rolling-Shutter Image). These distortions are not only visually unpleasing but also detrimental to downstream tasks [1, 2, 3, 4, 5]. Prior methods for RS removal can be categorized as multi-frame and single-frame based. The former relies on temporal motion in consecutive frames for motion compensation, while the latter solely depends on a single frame for restoration. Single-frame RS correction is particularly challenging yet crucial in data-scarce situations. Many existing methods, including non-learning-based approaches [6, 7, 8, 9, 10], rely on salient structures, such as straight lines. However, they may falter in cases where such salient structures are absent.
@@
The majority of deep RS correction methods are multi-frame based [11, 12, 13, 14, 15], with only a few considering a single frame as input. Among single-frame methods, Rengarajan et al. [7] directly learn the mapping between global shutter (GS) and RS frames, facing the challenge of large solution space for row-wise motion estimation. Zhuang et al. [9] addressed this by estimating an additional depth map, although depth estimation from RS frames is inherently ill-posed. Recently, Yan et al. [16] introduced a deep homography mixture model, achieving the current best performance by embedding motion in a subspace and learning coefficients to combine pre-learned motion flow bases.
@@
Recent advancements in generative models [17, 18, 19], including Diffusion Models (DMs) [20, 21], have significantly advanced the field of Artificial Intelligence. Notably, DMs excel in performing various motion-related tasks, notably in human motion generation [22] and in estimating depth/optical flow [23]. Meanwhile, DMs have shown strength in solving ill-posed problems, such as image restoration [24, 25, 26, 27], where inferring accurate solutions from ambiguous data is difficult. Therefore, drawing on these insights and inspirations, in this work, we introduce a novel single-frame RS correction method based on DMs, namely RS-Diffusion. DMs are adept at generating data from Gaussian distributions through multiple sampling steps and can be conditioned on additional information, such as text or images. Leveraging the transformative capabilities of DMs, we are capable of correcting RS images to GS images. Our framework, built on CFG [28] and DDIM [29], uses a downsampled RS image $I_{RS}$ as a conditioning element, concatenated with Gaussian noise. This input is processed by DMs in very few steps to produce a motion field $G$, which is then applied to remap $I_{RS}$ to eliminate RS artifacts. Additionally, we introduce a patch-attention module based on prior RS motion patterns to enhance results. We visualize the produced correcting field and our corrected result in the first row of Fig. 1.
@@
On the other hand, high-quality datasets play a crucial role. To meet the demand, we adhere to two essential criteria [30]: the label criterion, which requires precise alignment of RS correction motion as ground-truth labels between RS-GS pairs, and the realism criterion, which ensures that both the image content and RS motion remain realistic. However, existing RS datasets often fall short of meeting these criteria. The synthesized dataset can provide accurate labels, but violates realism criterion. In contrast, the captured images can suffice the realism, but the label is not accurate. For example, capturing with two cameras, one RS and one GS, suffers from parallax between two cameras.
@@
To resolve these issues, we introduce Intra Gyro Field (IGF) pipeline, leveraging Inertial Measurement Unit (IMU) gyroscope sensors to record camera rotations during capture. Specifically, we firstly achieve accurate frame-gyro data synchronization. Next, these rotations are translated into a series of homography matrices, further converted into motion fields as RS correction ground-truth (GT) labels. With the help of IGF, we create the RS-Real dataset, which, for the first time, satisfies both the label and realism criteria in RS correction. The dataset contains 44,000 train and 1,000 test samples. A case is shown in Fig. 1, where the second row shows a captured RS image, the IGF between RS-GS pair and its GS image.
@@
In summary, our diffusion model-based framework advances the state-of-the-art in single-image RS correction, while could run inference in real-time speed, i.e., up to 28.1 ms per frame on one NVIDIA 2080Ti. The RS-Real dataset, containing high-quality training pairs, addresses the scarcity of qualified datasets in the RS task. Our contributions include:
@@
  • We introduce a pipeline for correcting RS images with recorded IMU gyro readings, which delivers accurate rectified RS images as ground-truth labels for training and testing, yielding the first real RS dataset, RS-Real.
@@
Figure 1: Illustration of our results and the proposed dataset: the first row demonstrates that for a rolling-shutter (RS) image, our proposed RS-Diffusion framework produces a correcting field (in the optical flow visualization, different colors represent different directions and intensities of motion) leading to the corresponding corrected RS image. The second row illustrates our proposed RS-Real dataset, which includes, from left to right, a different RS image taken in real-world scenes, the ground-truth correcting field named Intra Gyro Field (IGF), which is used to correct the RS image to a global-shutter (GS) image, and the corresponding ground-truth GS image.
II Related Work
II-A Rolling Shutter Correction Methods.
Multi-frame rectification methods use spatial-temporal information, with investigations into per-pixel motion vectors [11], image feature tracking [12], and affine models [31], now enhanced by 3D data [13]. Advances in deep learning include pixel-wise velocity estimation [32], the use of optical flow for deep feature warping [33], adaptive warping techniques [34], transformer-based one-stage networks [35], and studies on bilateral motion fields [36], with recent work focusing on RS removal and deblurring [37]. Additionally, some studies explore the restoration of GS images by leveraging dual-RS cameras [38] and then achieve temporal super-resolution [39] via a self-supervised learning strategy [40]. Classical single-image methods mostly utilize salient lines for RS correction [6, 41]. Deep learning for single RS image correction is less common but includes neural networks trained on paired images [7], employing motion models and depth map estimations [9]. A novel deep homography mixture model shows state-of-the-art results [16]. We introduce diffusion models (DMs) for single RS rectification, demonstrating their effectiveness and generalizability in motion extraction.
II-B Diffusion Models.
Diffusion models (DMs), based on stochastic diffusion processes [42], efficiently transform data distributions through Gaussian transitions and iterative denoising using the data distribution gradient [43]. DDPM [20] employs discrete steps in this process, and Song et al. [21] further refines the methodology via stochastic differential equations (SDE). DDIM [29] accelerates reverse sampling through subsequence sampling and ordinary differential equations (ODEs). Conditioned data generation progresses with classifier-based [44, 45] and classifier-free (CFG) techniques [28], including cross-attention-based integration [46] and distinguishing between range-null space for conditioned generation [47]. DMs also contribute to motion-centric tasks, from video generation [48], image rectangling [49], hand mesh rendering [50] and homography rendering/estimation [51], optical flow [23] to human motion [22]. Our work extends DMs to the domain of rolling shutter removal.
II-C Gyroscope-based Motion Methods
Gyroscopes are pivotal for diverse applications such as video stabilization [52, 53], deblurring [54, 55], optical image stabilization (OIS) [56], simultaneous localization and mapping  [57], ego-motion [58, 59], and human pose and motion estimation [60, 61], with growing using in RS correction if gyroscope data is accessible during the capturing [62]. In this work, instead of directly using online captured gyroscope data to promote the performances of different applications. We use the gyroscope data to create a dataset. Our work utilizes gyroscope synchronization at the Android HAL for precise calibration [63, 59], enabling the creation of a RS-Real dataset for rolling shutter effect research, incorporating authentic RS pairings for enhanced training and validation. Note that, our method does not need gyro readings during the model inference.
@@ II-C Gyroscope-based Motion Methods
III Method
III-A Overview
We present a pipeline named Intra Gyro Field (IGF) designed to create a realistic and precisely annotated dataset through combined engineering and algorithmic efforts, as discussed in Sec. III-B. Furthermore, we propose a novel and general framework based on diffusion models, incorporating a specialized patch attention (PA) module, which is aimed at correcting rolling shutter (RS) images to global shutter (GS) quality, as described in Sec. III-C.
III-B Intra Gyro Field
Gyroscope records the relative camera 3D rotation in a relative time and it is possible to convert the gyroscope readings into a 2D motion field to achieve alignment between two consecutive frames [64]. In this chapter, we demonstrate how to leverage the gyroscope to correct RS effect for a single image and thus propose an Intra Gyro Field (IGF). Specifically, given the 3-axis angular velocity readings (row $\mathbf{v_{r}}$, pitch $\mathbf{v_{p}}$ and yaw $\mathbf{v_{y}}$) and relative time intervals $\Delta t$ between gyro readings. The 3-axis angular angles ($\mathbf{\angle r,\angle p,\angle y}$) can be computed as:
@@ III-B Intra Gyro Field
Then we utilize the Rodrigues Formula [65] to convert rotation angles into the rotation matrix $\mathbf{R}(\Delta t)\in SO(3)$. In order to bridge the relationship between 3D camera pose and 2D image motion, we choose to use the homography matrix $\mathbf{H}$. The homography matrix represents the relationship between two different perspectives of the same scene, more specifically, the conditions for its validity are satisfied when the camera motion is purely rotational or when the contents lie in the same plane [66]. Theoretically, given 3D rotation matrix $\mathbf{R}(\Delta t)$ and translation vector $\mathbf{t}(\Delta t)$), we can represent $\mathbf{H}$ as:
@@
where $\mathbf{{n}}^{T}$ represents the normal vector of a certain plane, $d$ is the distance between the camera center and the plane, and $\mathbf{K}$ is the camera intrinsic matrix.
In this work, we do not incorporate translations and only use gyroscope data to model the homography matrix, the reasons are threefold. Firstly, the rolling shutter effect caused by camera shake occurs primarily due to rotational movements [52]. Secondly, even though translations can be gathered from accelerometer data, they are significantly less accurate than rotational measurements [67, 68]. Lastly, according to Eq. 2, we can find that translation is correlated to depth, but accurately estimating depth maps can be another non-trivial problem. As a result, we can formulate a rotational-only homography:
@@
The gyroscope records $N$ readings during the capture of one image, it thus can record and produce $N$ camera rotation matrices ($\mathbf{R}_{1}(\Delta t),\mathbf{R}_{2}(\Delta t),\cdots,\mathbf{R}_{N}(\Delta t)$). If we split the image from top to bottom into $N$ patches ($\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{N}$), these rotation matrices correspond to the motion between the first row and bottom row of patches as illustrated in Fig. 2. Moreover, the inter-patch motion is approximately to be smooth [56, 69], so we can apply the spherical linear interpolation (SLERP) to interpolate the motion to avoid the discontinuities across row patches. Subsequently, we follow Eq. 3 to produce $N$ homography matrices ($\mathbf{H}_{1},\mathbf{H}_{2},\cdots,\mathbf{H}_{N}$) to model the 2D relationship between patches. To this end, we follow previous method [70] to convert the array of homography matrix into a motion field, it is achieved by transforming grid points by their corresponding homography and subtracting them. Lastly, this motion field is IGF, denoted as $\mathbf{G}\in\mathbf{R}^{2\times H\times W}$, which is capable of correcting the RS effect within an image:
@@
Figure 3: RS-Real dataset contains data pairs featuring various RS motion patterns and different intensities, all of which are captured in diverse scenes.
Dataset Collecting. Our proposed IGF effectively addresses the challenges found in previous datasets, such as unrealistic data distributions and spatio-temporal synchronization issues. By recording the RS effect with a gyroscope and subsequently recovering from it, our approach circumvents synchronization challenges by using only one camera at a time. As a result, we can flexibly capture a diverse range of realistic scenes, spanning various indoor and outdoor environments, and subjecting the camera to multiple motion patterns and speeds. This method enables us to gather a comprehensive dataset of realistic RS images, denoted as $\mathbf{X}_{RS}=\left\{\mathbf{I}_{RS}^{0},\mathbf{I}_{RS}^{1},\ldots,\mathbf{I}_{RS}^{k}\right\}$, along with their corresponding IGFs, $\mathbf{X}_{IGF}=\left\{\mathbf{G}^{0},\mathbf{G}^{1},\ldots,\mathbf{G}^{k}\right\}$. After applying RS removal using Eq. 4, we obtain the set of GS images, $\mathbf{X}_{GS}=\left\{\mathbf{I}_{GS}^{0},\mathbf{I}_{GS}^{1},\ldots,\mathbf{I}_{GS}^{k}\right\}$. We refer to this novel dataset as RS-Real. Examples from the dataset are illustrated in Fig. 3.
III-C Diffusion Models
To correct distortions resulting from the RS effect in general cases, we utilize diffusion models adept at managing the distributional transformations of images captured under both rolling shutter (RS) and global shutter (GS) conditions.
Specifically, we leverage the diffusion models as proposed by Sohl-Dickstein et al. [42] and Ho et al. [20], which define a Markov chain process over $T$ steps to incrementally introduce Gaussian noise into the original data distribution $\mathbf{x}_{0}\sim q(\mathbf{x})$. This stepwise infusion of noise generates a sequence of increasingly distorted states ${\mathbf{x}_{1},\ldots,\mathbf{x}_{T}}$, defined as forward diffusion, and is mathematically represented as:
@@ III-C Diffusion Models
This inversion transforms the signal from the noise-dominant Gaussian distribution back into the target data distribution. To refine this reconstruction process and improve the quality of the generated samples, we can integrate conditional variables $\mathbf{y}$ into the model as described by CFG [28]. These conditions are merged with the noisy data transitions in the model:
@@
III-D Rolling Shutter Removal Module
Our framework is built upon CFG [28] and DDIM [29]. We illustrate the pipeline in Fig. 4, we designate IGFs as the initial data distribution, $\mathbf{x}_{0}$, and the RS images $\mathbf{I}_{RS}$ as conditions $\mathbf{y}$. This configuration endows the model with an “image-to-motion” capability. During the training process, $\mathbf{x}_{0}$ is noised via forward diffusion to $\mathbf{x}_{t}$, then we concatenate the $\mathbf{I}_{RS}$, $\mathbf{x}_{t}$ and time embedding to fed into the U-net network $\theta$. It is imperative to note that this concatenation is introduced at various layers within the network. Finally, the output of $\theta$ is processed through the Patch-Attention module, resulting in the generated outputs $\hat{\mathbf{x}}_{0}$.
@@ III-D Rolling Shutter Removal Module
In addition to the MSE Loss, which facilitates learning the distribution transformation from RS images to their respective IGF, we further propose to constrain the network with an extra conditional loss. Specifically, we warp the original RS images $\mathbf{I}_{RS}$ via the computed IGF $\mathbf{\hat{x}}_{0}$ to produce corrected RS images $\mathbf{I}_{GS}^{\prime}$:

### v2

@@ Abstract
We present RS-Diffusion, the first Diffusion Models-based method for single-frame Rolling Shutter (RS) correction. RS artifacts compromise visual quality of frames due to the row-wise exposure of CMOS sensors. Most previous methods have focused on multi-frame approaches, using temporal information from consecutive frames for the motion rectification. However, few approaches address the more challenging but important single frame RS correction. In this work, we present an “image-to-motion” framework via diffusion techniques, with a designed patch-attention module. In addition, we present the RS-Real dataset, comprised of captured RS frames alongside their corresponding Global Shutter (GS) ground-truth pairs. The GS frames are corrected from the RS ones, guided by the corresponding Inertial Measurement Unit (IMU) gyroscope data acquired during capture. Experiments show that RS-Diffusion surpasses previous single-frame RS methods, demonstrates the potential of diffusion-based approaches, and provides a valuable dataset for further research.
Codehttps://github.com/lhaippp/RS-Diffusion
Datasetshttps://huggingface.co/Lhaippp/RS-Diffusion
Introduction
@@
Rolling shutter (RS) is a common effect encountered when capturing images with CMOS sensors. It results from varying exposure times for different rows in each frame, causing artifacts like distorted straight lines and skewed image content, as shown in Fig. 1 (RS Image). These distortions are not only visually unpleasing but also detrimental to downstream tasks (Hedborg et al. 2012; Saurer, Pollefeys, and Lee 2016; Albl, Kukelova, and Pajdla 2015; Saurer et al. 2013). Prior methods for RS removal can be categorized as multi-frame and single-frame based. The former relies on temporal motion in consecutive frames for motion compensation, while the latter solely depends on a single frame for restoration. Single-frame RS correction is particularly challenging yet crucial in data-scarce situations. Many existing methods, including non-learning-based approaches (Rengarajan, Rajagopalan, and Aravind 2016; Rengarajan, Balaji, and Rajagopalan 2017; Purkait, Zach, and Leonardis 2017; Zhuang et al. 2019; Kandula, Kumar, and Rajagopalan 2020), rely on salient structures, such as straight lines. However, they may falter in cases where such salient structures are absent.
Figure 1: Illustration of the proposed dataset and our results. The first row features a rolling-shutter (RS) image captured in realistic scenes, along with the corresponding ground-truth global-shutter (GS) image. The ground-truth flow used for correcting the RS image to the GS image is displayed on the left side of the second row. In the bottom right, we showcase our corrected RS image.
@@
The majority of deep RS correction methods are multi-frame based (Liang, Chang, and Chen 2008; Ringaby and Forssén 2012; Zhuang, Cheong, and Hee Lee 2017; Vasu, Rajagopalan et al. 2018), with only a few considering a single frame as input. Among single-frame methods, Rengarajan et al. (Rengarajan, Balaji, and Rajagopalan 2017) directly learn the mapping between global shutter (GS) and RS frames, facing the challenge of large solution space for row-wise motion estimation. Zhuang et al. (Zhuang et al. 2019) addressed this by estimating an additional depth map, although depth estimation from RS frames is inherently ill-posed. Recently, Yan et al. (Yan et al. 2023) introduced a deep homography mixture model, achieving the current best performance by embedding motion in a subspace and learning coefficients to combine pre-learned motion flow bases.
@@
Recent advancements in generative models (Brown et al. 2020; Goodfellow et al. 2020; Ouyang et al. 2022), including Diffusion Models (DMs) (Ho, Jain, and Abbeel 2020; Song et al. 2020), have significantly advanced the field of Artificial Intelligence. Notably, DMs excel in performing various motion-related tasks, notably in human motion generation (Tevet et al. 2022), in estimating depth/optical flow (Saxena et al. 2023), and in homography data rendering (Li et al. 2024). Meanwhile, DMs have shown strength in solving ill-posed problems, such as image restoration (Gao et al. 2023; Yinhuai, Jiwen, and Jian 2022), where inferring accurate solutions from ambiguous data is difficult. Therefore, drawing on these insights and inspirations, in this work, we introduce a novel single-frame RS correction method based on DMs, namely RS-Diffusion. DMs are adept at generating data from Gaussian distributions through multiple sampling steps and can be conditioned on additional information, such as images. Leveraging the transformative capabilities, we are capable of correcting RS images to GS images. Our framework, built on CFG (Ho and Salimans 2022) and DDIM (Song, Meng, and Ermon 2020), uses a downsampled RS image $I_{RS}$ as a conditioning element, concatenated with Gaussian noise. This input is processed by DMs in very few steps to produce a motion field $G$, which is then applied to remap $I_{RS}$ to eliminate RS artifacts. Additionally, we introduce a patch-attention module based on prior RS motion patterns to enhance results. We visualize our corrected result on the right side of the second row in Fig. 1.
@@
On the other hand, high-quality datasets play a crucial role. To meet the demand, we adhere to two essential criteria (Han et al. 2022): the label criterion, which requires precise alignment of the RS correction motion as ground-truth labels between RS-GS pairs, and the realism criterion, which ensures that both the image content and RS motion remain realistic. However, existing RS datasets often fall short of meeting these criteria. The synthesized dataset can provide accurate labels but violates realism criterion. In contrast, the captured images can suffice the realism, but the label is not accurate.
@@
To resolve these issues, we introduce Intra Gyro Field (IGF) pipeline, leveraging Inertial Measurement Unit (IMU) gyroscope sensors to record camera rotations during capture. Specifically, we firstly achieve accurate frame-gyro data synchronization. Next, these rotations are translated into a series of homography matrices, further converted into motion fields as RS correction ground-truth (GT) labels. With the help of IGF, we create the RS-Real dataset, which satisfies both the label and the realism criteria in RS correction. The dataset contains 40,000 train and 1,000 test samples. A case is shown in Fig. 1, which shows a captured RS image, the IGF between the RS-GS pair and its GS image.
@@
In summary, our diffusion model-based framework advances the state-of-the-art in single-image RS correction, while it could run inference in real-time speed, i.e., up to 28.1 ms per frame on one NVIDIA 2080Ti. The RS-Real dataset, containing high-quality training pairs, addresses the scarcity of qualified datasets in the RS task. Our contributions include:
@@
  • We introduce a pipeline for correcting RS images with recorded IMU gyro readings, which delivers accurate rectified RS images as ground-truth labels for training and testing, yielding a realistic RS dataset, RS-Real.
@@
Related Works
Rolling Shutter Correction Methods
Multi-frame rectification methods use spatial-temporal information, with investigations into per-pixel motion vectors (Liang, Chang, and Chen 2008), image feature tracking (Ringaby and Forssén 2012), and affine models (Baker et al. 2010), now enhanced by 3D data (Zhuang, Cheong, and Hee Lee 2017). Deep learning advances include pixel-wise velocity estimation (Liu et al. 2020) and optical flow for deep feature warping (Fan, Dai, and He 2021). Classical single-image methods mostly utilize salient lines for RS correction (Rengarajan, Rajagopalan, and Aravind 2016). Deep learning for single RS image correction is less common but includes neural networks trained on paired images (Rengarajan, Balaji, and Rajagopalan 2017), employing motion models and depth map estimations (Zhuang et al. 2019). A novel deep homography mixture model shows state-of-the-art results (Yan et al. 2023). We introduce diffusion models (DMs) for single RS rectification, demonstrating their effectiveness and generalizability in motion extraction.
Diffusion Models
Diffusion models (DMs), based on stochastic diffusion processes (Sohl-Dickstein et al. 2015), efficiently transform data distributions through Gaussian transitions and iterative denoising using the data distribution gradient (Song and Ermon 2019). DDPM (Ho, Jain, and Abbeel 2020) employs discrete steps in this process, and Song et al. (Song et al. 2020) further refines the methodology via stochastic differential equations (SDE). DDIM (Song, Meng, and Ermon 2020) accelerates reverse sampling through subsequence sampling and ordinary differential equations (ODEs). Conditioned data generation progresses with classifier-based (Dhariwal and Nichol 2021; Liu et al. 2023) and classifier-free (CFG) techniques (Ho and Salimans 2022). DMs also contribute to motion-centric tasks, from video generation (Ni et al. 2023), video frame interpolation (Danier, Zhang, and Bull 2024), object tracking (Luo et al. 2024) to Molecule generation (Huang et al. 2023). Our pioneering work extends DMs to the domain of rolling shutter removal.
@@
Gyroscope-based Motion Methods
Gyroscopes are pivotal for diverse applications such as video stabilization (Karpenko et al. 2011; Bell, Troccoli, and Pulli 2014), optical image stabilization (OIS) (Liu et al. 2021), estimating homography/optical flow (Li, Luo, and Liu 2021; Li et al. 2023), and simultaneous localization and mapping (SLAM) (Huang and Liu 2018), with growing using in RS correction if gyroscope data is accessible during the capturing (Mo, Islam, and Sattar 2022). In this work, instead of directly using online captured gyroscope data to promote the performances of different applications. We use the gyroscope data to create a dataset. Our work utilizes gyroscope synchronization at the Android HAL for precise calibration (Jia and Evans 2013; Bloesch et al. 2014), enabling the creation of a RS-Real dataset for rolling shutter research, incorporating authentic RS pairings for enhanced training and validation. Note that, our method doesn’t need gyro readings during the model inference.
Method
Overview
We present a pipeline named Intra Gyro Field (IGF) designed to create a realistic and precisely annotated dataset through combined engineering and algorithmic efforts, as discussed in Section Intra Gyro Field. Furthermore, we propose a novel and general framework based on diffusion models, incorporating a specialized patch attention (PA) module, which is aimed at correcting rolling shutter (RS) images to global shutter (GS) quality, as described in Section Diffusion Models.
Intra Gyro Field
Gyroscope records the relative camera 3D rotation in a relative time and it is possible to convert the gyroscope readings into a 2D motion field to achieve alignment between two consecutive frames (Li et al. 2023). In this chapter, we demonstrate how to leverage the gyroscope to correct RS effect for a single image and thus propose an Intra Gyro Field (IGF). Specifically, given the 3-axis angular velocity readings (row $\mathbf{v_{r}}$, pitch $\mathbf{v_{p}}$ and yaw $\mathbf{v_{y}}$) and relative time intervals $\Delta t$ between gyro readings. The 3-axis angular angles ($\mathbf{\angle r,\angle p,\angle y}$) can be computed as:
@@
Then we utilize the Rodrigues Formula to convert rotation angles into the rotation matrix $\mathbf{R}(\Delta t)\in SO(3)$. In order to bridge the relationship between 3D camera pose and 2D image motion, we choose to use the homography matrix $\mathbf{H}$. The homography matrix represents the relationship between two different perspectives of the same scene, more specifically, the conditions for its validity are satisfied when the camera motion is purely rotational or when the contents lie in the same plane (Hartley and Zisserman 2003). Theoretically, given 3D rotation matrix $\mathbf{R}(\Delta t)$ and translation vector $\mathbf{t}(\Delta t)$), we can represent $\mathbf{H}$ as:
@@
where $\mathbf{{n}}^{T}$ represents the normal vector of a certain plane, $d$ is the distance between the camera center and the plane, while $\mathbf{K}$ is the camera intrinsic matrix.
Figure 4: A glance at the RS-Real dataset reveals that it contains data pairs featuring various RS motion patterns and different intensities, all of which are captured in diverse scenes.
In this work, we do not incorporate translations and only use gyroscope data to model the homography matrix, the reasons are threefold. Firstly, the rolling shutter effect caused by camera shake occurs primarily due to rotational movements (Karpenko et al. 2011). Secondly, even though translations can be gathered from accelerometer data, they are significantly less accurate than rotational measurements (Forssén and Ringaby 2010; Joshi et al. 2010). Lastly, according to Eq. 2, we can find that translation is correlated to depth, but accurately estimating depth maps can be another non-trivial problem. As a result, we can formulate a rotational-only homography:
@@
The gyroscope records $N$ readings during the capture of one image, it thus can produce $N$ camera rotation matrices ($\mathbf{R}_{1}(\Delta t),\mathbf{R}_{2}(\Delta t),\cdots,\mathbf{R}_{N}(\Delta t)$). If we split the image from top to bottom into $N$ patches ($\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{N}$), these rotation matrices correspond to the motion between the first row and bottom row of patches as illustrated in Fig. 2. Moreover, the inter-patch motion is approximately to be smooth (Liu et al. 2021; Dai, Li, and Kneip 2016), so we can apply the spherical linear interpolation (SLERP) to interpolate the motion to avoid the discontinuities across row patches. Subsequently, we follow Eq. 3 to produce $N$ homography matrices ($\mathbf{H}_{1},\mathbf{H}_{2},\cdots,\mathbf{H}_{N}$) to model the 2D relationship between patches. To this end, we follow previous method (Li, Luo, and Liu 2021) to convert the array of homography matrix into a motion field, it is achieved by transforming grid points by their corresponding homography and subtracting them. Lastly, this motion field is IGF, denoted as $\mathbf{G}\in\mathbf{R}^{2\times H\times W}$, which is capable of correcting the RS effect within an image:
@@
Dataset Collecting. Our proposed IGF effectively addresses the challenges found in previous datasets, such as unrealistic data distributions and spatio-temporal synchronization issues. By recording the RS effect with a gyroscope and subsequently recovering from it, our approach circumvents synchronization challenges by using only one camera at a time. As a result, we can flexibly capture a diverse range of realistic scenes, spanning various indoor and outdoor environments, and subjecting the camera to multiple motion patterns and speeds. This method enables us to gather a comprehensive dataset of realistic RS images, denoted as $\mathbf{X}_{RS}=\left\{\mathbf{I}_{RS}^{0},\mathbf{I}_{RS}^{1},\ldots,\mathbf{I}_{RS}^{k}\right\}$, along with their corresponding IGFs, $\mathbf{X}_{IGF}=\left\{\mathbf{G}^{0},\mathbf{G}^{1},\ldots,\mathbf{G}^{k}\right\}$. After applying RS removal using Eq. 4, we obtain the set of GS images, $\mathbf{X}_{GS}=\left\{\mathbf{I}_{GS}^{0},\mathbf{I}_{GS}^{1},\ldots,\mathbf{I}_{GS}^{k}\right\}$. We refer to this dataset as RS-Real. Examples from the dataset are illustrated in Fig. 4.
Diffusion Models
To correct distortions resulting from the rolling shutter effect in general cases, we utilize diffusion models adept at managing the distributional transformations of images captured under both RS and GS conditions.
Specifically, we leverage the diffusion models as proposed by Sohl-Dickstein et al. (Sohl-Dickstein et al. 2015) and Ho et al. (Ho, Jain, and Abbeel 2020), which define a Markov chain process over $T$ steps to incrementally introduce Gaussian noise into the original data distribution $\mathbf{x}_{0}\sim q(\mathbf{x})$. This stepwise infusion of noise generates a sequence of increasingly distorted states ${\mathbf{x}_{1},\ldots,\mathbf{x}_{T}}$, defined as forward diffusion, and is mathematically represented as:
@@
This inversion transforms the signal from the noise-dominant Gaussian distribution back into the target data distribution. To refine this reconstruction process and improve the quality of the generated samples, we can integrate conditional variables $\mathbf{y}$ into the model as described by CFG (Ho and Salimans 2022). These conditions are merged with the noisy data transitions in the model:
@@
Rolling Shutter Removal Module
Our framework is built upon CFG (Ho and Salimans 2022) and DDIM (Song, Meng, and Ermon 2020). We illustrate the pipeline in the Fig. 3, we designate IGFs as the initial data distribution, $\mathbf{x}_{0}$, and the RS images $\mathbf{I}_{RS}$ as conditions $\mathbf{y}$. This configuration endows the model with an “image-to-motion” capability. During the training process, $\mathbf{x}_{0}$ is noised via forward diffusion to $\mathbf{x}_{t}$, then we concatenate the $\mathbf{I}_{RS}$, $\mathbf{x}_{t}$ and time embedding to fed into the U-net network $\theta$. It is imperative to mention that this concatenation is introduced at various layers within the network. Finally, the output of $\theta$ is processed through the Patch-Attention module, resulting in the generated outputs $\hat{\mathbf{x}}_{0}$.
@@
In addition to the standard MSE Loss, which facilitates learning the distribution transformation from RS images to their respective IGF, we further propose to constrain the network with an extra conditional loss. Specifically, we warp the original RS images $\mathbf{I}_{RS}$ via the computed IGF $\mathbf{\hat{x}}_{0}$ to produce corrected RS images $\mathbf{I}_{GS}^{\prime}$: